# Think, Act, Learn: A Framework for Autonomous Robotic Agents using Closed-Loop Large Language Models

### v1

@@ Abstract
Index Terms:
Embodied AI, Large Language Models, Robotics, Autonomous Learning, Task Planning, Feedback Control, Human-Robot Interaction.
@@ I Introduction
The advent of Large Language Models (LLMs) has marked a pivotal moment in artificial intelligence, demonstrating remarkable capabilities in understanding, generating, and reasoning with human language. This progress has catalyzed a paradigm shift from passive text generation to active environmental interaction, giving rise to the concept of autonomous agents [1]. These agents, powered by foundation models, are envisioned to perform complex tasks on behalf of users, fundamentally altering the landscape of human-computer interaction. The initial proving ground for this technology has been the World Wide Web, a vast and semi-structured environment ripe for automation. Pioneering works have introduced sophisticated benchmarks and datasets like Mind2Web [2] and WebArena [3], alongside innovative agents such as WebVoyager [4], which aim to create generalist agents for the web. These efforts have highlighted the potential of agents to tackle complex, real-world web tasks [5, 6] and have even suggested that for many tasks, the browser is all you need [7].
@@
Despite significant progress in web automation, the ultimate goal is to create generalist agents that can operate across any digital interface, including desktop and mobile applications. This leap presents a distinct set of challenges, moving from structured HTML to unstructured pixel-based graphical user interfaces (GUIs). Recent efforts have begun to tackle this frontier, with projects like AppAgent [8] and Mobile-Agent [9] demonstrating impressive capabilities on smartphones, supported by new evaluation platforms like Mobile-Env [10] and large-scale datasets [11]. Concurrently, the control of desktop environments is being explored through platforms like Desktop-Env [12] and pilot studies on specific operating systems [13], with a vision towards universal computer control [14]. This transition requires agents to handle a wider array of inputs, from traditional GUIs to more diverse interaction modalities. For instance, recognizing user intent can involve interpreting gestures, a task that has been explored using various sensing technologies, including attention-based WiFi recognition [15], and advances in screen-content understanding that help disambiguate on-screen intent [16]. Furthermore, ensuring robust interaction requires dealing with open-set conditions where unknown gestures may appear, a problem addressed by frameworks like WiOpen [17].
@@
At the heart of a generalist UI agent lie three fundamental capabilities: perception, reasoning, and action. The perception module, responsible for interpreting the visual interface, has been revolutionized by Vision-Language Models (VLMs) like Flamingo and the powerful GPT-4V [18]. Specialized models such as CogAgent [19] and ScreenAI [20] have been developed specifically for UI understanding, further enhanced by prompting techniques like Set-of-Mark to improve visual grounding [21]. This visual understanding often requires deep context, such as inferring relationships through graph structures [22] and leveraging multimodal dialogue understanding toolkits [23], or propagating information along event lines [24]. The reasoning and planning capabilities of agents are built upon seminal frameworks like Chain-of-Thought [25], which elicits reasoning, and ReAct [26], which synergizes reasoning with acting. These foundational ideas are being extended to allow models to teach themselves to use tools [27] and to reason algorithmically [28]. The action generation component must then translate these plans into concrete operations, a process that can be informed by understanding micro-actions [29] , benchmarked in instruction-following-with-vision settings [30] and employing efficient temporal filtering for dynamic scenes [31].
@@
The learning paradigm for such agents is a critical area of research. While early approaches relied on direct supervision from web-scale command data [32] or simple vision-to-action mappings [33], the field is rapidly moving towards more autonomous learning methods. Reinforcement Learning (RL) has long been a staple for teaching agents in interactive environments [34], and modern approaches reframe RL as a sequence modeling problem using Decision Transformers [35]. To improve learning efficiency and performance, hybrid methods are emerging. The Reflexion framework, for instance, enables agents to learn from past failures through verbal reinforcement learning [36], while other works explore self-improvement through human-in-the-loop play [37]. A significant challenge in learning from real-world data is the presence of noise; developing methods for reliable label noise suppression is therefore crucial for robust model training, particularly in tasks like facial expression recognition that inform affective computing . In parallel, broader analyses of decision-making underscore the importance of data quality and robustness in agent systems [1]. Moreover, as agents become more distributed, federated learning approaches that incorporate experience-driven model migration can optimize performance in heterogeneous edge networks.
@@
Building and evaluating these complex systems necessitates a diverse ecosystem of benchmarks and datasets. Beyond web-specific platforms, there is a growing need for benchmarks that test instruction-following with vision in challenging scenarios [30] and assess mathematical reasoning in visual contexts [38]. The community has developed comprehensive evaluation suites like AgentBench [39] and GAIA [40] to rigorously test LLMs as generalist assistants. The scope of agent interaction is also expanding beyond the visual domain. Research into WiFi sensing, for example, has demonstrated the potential for recognizing human activities with anti-interference methods [41], with emerging unified multi-sensor autonomy frameworks offering complementary perspectives [42]; it also enables unobtrusive emotion recognition by fusing WiFi with vision [43], and aligns with broader perspectives on generalist robotics [44]. Complementary simulation frameworks further emphasize multi-sensor autonomy [42]. Notably, commodity WiFi can even capture pulmonary function without contact , and reliability at the software level has been examined from a UI-testing perspective. This target-oriented sensing paradigm , together with the use of other RF technologies like RFID for activity recognition [45], points towards a future where agents perceive the world through multiple modalities. However, this expanded capability also introduces new security vulnerabilities, which must be addressed alongside rigorous software testing [46], such as attacks against PHY layer fingerprinting in WiFi authentication [47].
@@
This thesis introduces GUI-Learner, a novel framework for a general-purpose UI agent designed to learn to operate new software autonomously. Our work is situated at the intersection of multimodal foundation models, sequential decision-making, and autonomous learning. We draw inspiration from the progress in building generalist agents for both robotics [44] and complex virtual worlds [48], and we leverage insights from models tuned for in-context instruction following [49] and knowledge-based task solving [50]. The core challenge we address is enabling an agent to generalize from limited experience to unseen applications, a task that requires not only robust visual perception but also efficient learning from interaction. To this end, our model must be highly efficient, motivating research into techniques like multi-objective convex quantization for model compression [51]. We also contribute to the ongoing discussion about the capabilities of LLMs in software engineering tasks, such as automating UI-to-code generation [52] and UI-level testing [46]. Our primary contributions are: (1) A novel agent architecture that effectively grounds high-level instructions into low-level GUI actions. (2) A hybrid learning strategy combining imitation learning and offline reinforcement learning to achieve high sample efficiency and strong generalization. (3) A comprehensive evaluation on a diverse suite of desktop and web applications, demonstrating superior performance against existing methods and providing a blueprint for future foundation models for decision making [1]. Through this work, we aim to bridge the gap between current agent capabilities and the long-term vision of a truly universal AI assistant, capable of understanding and operating any software on demand. The ability to learn from dynamic visual information is paramount, building on foundational work in visual dialog [22], mathematical reasoning in visual contexts [38], and video grounding [31], and transferring knowledge from large-scale video datasets [53] within unified simulation frameworks [42] will be key to future progress. Finally, our work is benchmarked using standardized methodologies for evaluating visual question answering on screen content [16], ensuring our results are comparable and contribute to the broader scientific effort.
@@ II-A Foundation Models for Perception and Reasoning
The ability of an agent to perceive and understand a graphical user interface is paramount. Early efforts in GUI automation often relied on brittle, rule-based systems or direct access to the underlying view hierarchy, which lacked generalizability [33]. The recent success of agents is largely attributable to the power of foundation models. Vision-Language Models (VLMs) have become the de facto perception backbone. Seminal models like Flamingo demonstrated the potential of VLMs for few-shot learning in multimodal contexts , a crucial capability for adapting to new interfaces. The release of GPT-4V fundamentally changed the landscape, showcasing an unprecedented ability to reason about visual content directly from pixels [18]. This has spurred the development of specialized VLMs tailored for UI understanding, such as ScreenAI, which focuses on user experience evaluation [20], and CogAgent, which is optimized for fine-grained GUI element detection [19]. The effectiveness of these models can be further amplified by advanced prompting techniques like Set-of-Mark, which significantly improves visual grounding on crowded interfaces [21].
Beyond static perception, agents must understand dynamic scenes and dialog. This requires models that can perform efficient temporal filtering for tasks like video grounding [31] and understand context through mechanisms like graph inference and knowledge distillation [22] and multimodal dialogue understanding toolkits [23]. The ability to propagate information along event timelines, as explored in audio-visual research [24], is also conceptually relevant for tracking state changes in a GUI over time. Furthermore, a comprehensive understanding of on-screen content is often necessary, which is being addressed by dedicated benchmarks for visual question answering on screens [16] and multimodal dialogue understanding toolkits [23]. The reasoning capabilities of these agents are powered by the underlying LLM architecture, which has shown emergent abilities in planning and decomposition when guided by frameworks like Chain-of-Thought [25] and can be extended to perform complex algorithmic reasoning [28].
@@ II-B Autonomous Agents and Frameworks
An agent is more than just a perception model; it requires an agentic framework that enables it to plan and execute actions to achieve a goal. The ReAct framework was a landmark contribution, demonstrating how to effectively interleave reasoning steps with actions to create more robust and transparent agents [26]. This core idea has been extended in various ways, for example, by enabling agents to learn to use external tools and APIs autonomously [27]. The concept of self-improvement has also gained traction, with frameworks like Reflexion allowing agents to verbally reflect on past mistakes to improve future performance [36].
These agentic principles have been applied across different digital domains. In web automation, projects like WebVoyager [4] and others [5] have created agents capable of navigating complex websites. The scope has recently expanded to mobile devices with AppAgent [8] and Mobile-Agent [9], and to general desktop control with pilot studies on Windows [13] and ambitious projects aiming for universal computer control [14]. The ultimate vision is a generalist agent that can operate in any environment, a concept being pushed forward in complex domains like Minecraft [48] and robotics [44]. The knowledge required for such generalist agents can be substantial, leading to the development of open-source, knowledge-based agents that can solve complex tasks by querying external information sources [50]. Some approaches even aim to turn the agent into a ”programmer” that translates human instructions into executable code [54], blurring the lines between using and creating software.
@@ II-C Learning Paradigms for Interaction
Training these autonomous agents presents a significant research challenge. A primary methodology is imitation learning, or behavioral cloning, where the agent learns by mimicking expert demonstrations. This has been shown to be effective when large, web-scale datasets of commands are available [32]. However, imitation learning alone can suffer from compounding errors and may not generalize well to situations not seen in the training data. To overcome this, Reinforcement Learning (RL) is often employed, allowing agents to learn through trial and error. Early work applied RL to web navigation in richly-observed environments [34], and more recent paradigms like the Decision Transformer have elegantly reframed RL as a sequence modeling problem, making it highly compatible with modern LLM architectures [35].
Given the high cost of exploration in real-world applications, hybrid approaches are becoming increasingly popular. These methods often pre-train an agent with imitation learning and then fine-tune it with RL. Innovations in this area include self-improvement through human-in-the-loop interaction [37] and specialized fine-tuning strategies like Agent-Tuning to better align LLMs with agentic behaviors [55]. A critical and often-overlooked aspect of training is data quality. Real-world datasets are frequently noisy, and developing robust methods for label noise suppression is essential for training reliable models . Broader surveys of LLM-based agents also emphasize data quality and robustness . As these models grow in size, deploying them efficiently becomes a challenge, necessitating research into model compression techniques like multi-objective convex quantization [51]. For applications involving sensitive data, privacy-preserving training methods are required. Federated learning offers a promising solution, and recent work has focused on making it practical for heterogeneous edge networks through strategies like experience-driven model migration .
@@ II-D Sensing and Interaction Beyond the GUI
While our work focuses on visual GUI interaction, the broader field of human-computer interaction is exploring a rich tapestry of sensing modalities that provide valuable context for future generalist agents. For instance, WiFi-based sensing has been used for attention-based gesture recognition [15], and this line of work is complementary to advances in screen-content understanding for instruction following [16]. Robustness to open-set conditions—where unknown gestures may be present—has been investigated by frameworks like WiOpen [17] , with complementary insights from general-assistant evaluations [40] , and is related to benchmarking finer-grained behaviors such as micro-actions [29]. WiFi signals can also be used for human activity recognition, with techniques developed to mitigate co-channel interference [41], alongside broader evaluations of general AI assistants that stress realistic task settings [40]. The fusion of WiFi and vision further enables unobtrusive emotion recognition through combined analysis of gestures and facial expressions [43], which dovetails with efforts toward unified, multi-sensor autonomy [42]. In healthcare, non-contact pulmonary function monitoring with commodity WiFi has been demonstrated , and complementary perspectives on decision-making pipelines emphasize reliability [1]. This target-oriented sensing paradigm aligns with insights from generalist robotics [44]; beyond WiFi, other radio-frequency technologies like RFID can recognize fine-grained activities such as writing [45]. However, the reliance on physical-layer information opens security attack surfaces—e.g., undetectable attacks on PHY-layer fingerprinting for WiFi authentication [42]—underscoring the need for rigorous testing [46].
@@
The fusion of multiple modalities is a particularly promising frontier; for example, combining WiFi and vision has enabled unobtrusive emotion recognition through joint analysis of gestures and facial expressions [43]. This mirrors the architectural principles of powerful vision-language models like Flamingo, which are designed to process and reason about multimodal inputs [56, 57]. The application of these sensing technologies extends even to personal healthcare, where target-oriented sensing methods demonstrate significant potential . This paradigm of non-contact monitoring is a key application area for future foundation models for decision making [1], with specific systems successfully capturing pulmonary function using commodity WiFi [58]. In parallel, unified multi-sensor autonomy initiatives highlight cross-modal deployment challenges [42]. Furthermore, other radio-frequency technologies like RFID are being leveraged to recognize fine-grained activities such as writing [45]. This multi-sensor vision for autonomy, pursued in frameworks like UniSim , also introduces new security considerations. The reliance on physical layer information opens up novel attack surfaces, necessitating research into defending against threats like undetectable attacks on WiFi authentication systems [47].

### v2

@@ I Introduction
The advent of Large Language Models (LLMs) has marked a pivotal moment in artificial intelligence, demonstrating remarkable capabilities in understanding, generating, and reasoning with human language. This progress has catalyzed a paradigm shift from passive text generation to active environmental interaction, giving rise to the concept of autonomous agents [fu2024foundation]. These agents, powered by foundation models, are envisioned to perform complex tasks on behalf of users, fundamentally altering the landscape of human-computer interaction. The initial proving ground for this technology has been the World Wide Web, a vast and semi-structured environment ripe for automation. Pioneering works have introduced sophisticated benchmarks and datasets like Mind2Web [deng2023mind2web] and WebArena [zhou2023webarena], alongside innovative agents such as WebVoyager [he2024webvoyager], which aim to create generalist agents for the web. These efforts have highlighted the potential of agents to tackle complex, real-world web tasks [gur2023real, huang2024empirical] and have even suggested that for many tasks, the browser is all you need [podolny2024browser].
@@
Despite significant progress in web automation, the ultimate goal is to create generalist agents that can operate across any digital interface, including desktop and mobile applications. This leap presents a distinct set of challenges, moving from structured HTML to unstructured pixel-based graphical user interfaces (GUIs). Recent efforts have begun to tackle this frontier, with projects like AppAgent [yang2023appagent] and Mobile-Agent [fu2024mobile] demonstrating impressive capabilities on smartphones, supported by new evaluation platforms like Mobile-Env [zhang2024mobile] and large-scale datasets [rawles2023android]. Concurrently, the control of desktop environments is being explored through platforms like Desktop-Env [zhang2024desktopenv] and pilot studies on specific operating systems [cheng2024windows], with a vision towards universal computer control [cradle2024]. This transition requires agents to handle a wider array of inputs, from traditional GUIs to more diverse interaction modalities. For instance, recognizing user intent can involve interpreting gestures, a task that has been explored using various sensing technologies, including attention-based WiFi recognition [10090421], and advances in screen-content understanding that help disambiguate on-screen intent [li2024screenvqa]. Furthermore, ensuring robust interaction requires dealing with open-set conditions where unknown gestures may appear, a problem addressed by frameworks like WiOpen [10899398].
@@
At the heart of a generalist UI agent lie three fundamental capabilities: perception, reasoning, and action. The perception module, responsible for interpreting the visual interface, has been revolutionized by Vision-Language Models (VLMs) like Flamingo and the powerful GPT-4V [openai2023gpt4v]. Specialized models such as CogAgent [hong2023cogagent] and ScreenAI [baechle2024screenai] have been developed specifically for UI understanding, further enhanced by prompting techniques like Set-of-Mark to improve visual grounding [yang2023set]. This visual understanding often requires deep context, such as inferring relationships through graph structures and leveraging multimodal dialogue understanding toolkits [li2023v], or propagating information along event lines . The reasoning and planning capabilities of agents are built upon seminal frameworks like Chain-of-Thought [wei2022chain], which elicits reasoning, and ReAct [yao2023react], which synergizes reasoning with acting. These foundational ideas are being extended to allow models to teach themselves to use tools and to reason algorithmically [zelikman2024parsel]. The action generation component must then translate these plans into concrete operations, a process that can be informed by understanding micro-actions , benchmarked in instruction-following-with-vision settings [biswas2024avis] and employing efficient temporal filtering for dynamic scenes .
@@
The learning paradigm for such agents is a critical area of research. While early approaches relied on direct supervision from web-scale command data [li2023web] or simple vision-to-action mappings [sharma2020pix2act], the field is rapidly moving towards more autonomous learning methods. Reinforcement Learning (RL) has long been a staple for teaching agents in interactive environments [jia2019reinforcement], and modern approaches reframe RL as a sequence modeling problem using Decision Transformers [chen2021decision]. To improve learning efficiency and performance, hybrid methods are emerging. The Reflexion framework, for instance, enables agents to learn from past failures through verbal reinforcement learning [shinn2023reflexion], while other works explore self-improvement through human-in-the-loop play [zhang2024self]. A significant challenge in learning from real-world data is the presence of noise; developing methods for reliable label noise suppression is therefore crucial for robust model training, particularly in tasks like facial expression recognition that inform affective computing . In parallel, broader analyses of decision-making underscore the importance of data quality and robustness in agent systems [fu2024foundation]. Moreover, as agents become more distributed, federated learning approaches that incorporate experience-driven model migration can optimize performance in heterogeneous edge networks.
@@
Building and evaluating these complex systems necessitates a diverse ecosystem of benchmarks and datasets. Beyond web-specific platforms, there is a growing need for benchmarks that test instruction-following with vision in challenging scenarios [biswas2024avis] and assess mathematical reasoning in visual contexts [lu2024mathvista]. The community has developed comprehensive evaluation suites like AgentBench [liu2023agentbench] and GAIA to rigorously test LLMs as generalist assistants. The scope of agent interaction is also expanding beyond the visual domain. Research into WiFi sensing, for example, has demonstrated the potential for recognizing human activities with anti-interference methods [9075376], with emerging unified multi-sensor autonomy frameworks offering complementary perspectives [wang2023unisim]; it also enables unobtrusive emotion recognition by fusing WiFi with vision [10149418], and aligns with broader perspectives on generalist robotics [yuan2024towards]. Complementary simulation frameworks further emphasize multi-sensor autonomy [wang2023unisim]. Notably, commodity WiFi can even capture pulmonary function without contact , and reliability at the software level has been examined from a UI-testing perspective. This target-oriented sensing paradigm , together with the use of other RF technologies like RFID for activity recognition [feng2025rf], points towards a future where agents perceive the world through multiple modalities. However, this expanded capability also introduces new security vulnerabilities, which must be addressed alongside rigorous software testing [araujo2024can], such as attacks against PHY layer fingerprinting in WiFi authentication [10339891].
@@
This thesis introduces GUI-Learner, a novel framework for a general-purpose UI agent designed to learn to operate new software autonomously. Our work is situated at the intersection of multimodal foundation models, sequential decision-making, and autonomous learning. We draw inspiration from the progress in building generalist agents for both robotics [yuan2024towards] and complex virtual worlds [yuan2024ghost], and we leverage insights from models tuned for in-context instruction following [li2023otter] and knowledge-based task solving [zhou2024ok]. The core challenge we address is enabling an agent to generalize from limited experience to unseen applications, a task that requires not only robust visual perception but also efficient learning from interaction. To this end, our model must be highly efficient, motivating research into techniques like multi-objective convex quantization for model compression . We also contribute to the ongoing discussion about the capabilities of LLMs in software engineering tasks, such as automating UI-to-code generation [daza2023automating] and UI-level testing [araujo2024can]. Our primary contributions are: (1) A novel agent architecture that effectively grounds high-level instructions into low-level GUI actions. (2) A hybrid learning strategy combining imitation learning and offline reinforcement learning to achieve high sample efficiency and strong generalization. (3) A comprehensive evaluation on a diverse suite of desktop and web applications, demonstrating superior performance against existing methods and providing a blueprint for future foundation models for decision making [fu2024foundation]. Through this work, we aim to bridge the gap between current agent capabilities and the long-term vision of a truly universal AI assistant, capable of understanding and operating any software on demand. The ability to learn from dynamic visual information is paramount, building on foundational work in visual dialog , mathematical reasoning in visual contexts [lu2024mathvista], and video grounding , and transferring knowledge from large-scale video datasets [brohan2023rt] within unified simulation frameworks [wang2023unisim] will be key to future progress. Finally, our work is benchmarked using standardized methodologies for evaluating visual question answering on screen content [li2024screenvqa], ensuring our results are comparable and contribute to the broader scientific effort.
@@ II-A Foundation Models for Perception and Reasoning
The ability of an agent to perceive and understand a graphical user interface is paramount. Early efforts in GUI automation often relied on brittle, rule-based systems or direct access to the underlying view hierarchy, which lacked generalizability [sharma2020pix2act]. The recent success of agents is largely attributable to the power of foundation models. Vision-Language Models (VLMs) have become the de facto perception backbone. Seminal models like Flamingo demonstrated the potential of VLMs for few-shot learning in multimodal contexts , a crucial capability for adapting to new interfaces. The release of GPT-4V fundamentally changed the landscape, showcasing an unprecedented ability to reason about visual content directly from pixels [openai2023gpt4v]. This has spurred the development of specialized VLMs tailored for UI understanding, such as ScreenAI, which focuses on user experience evaluation [baechle2024screenai], and CogAgent, which is optimized for fine-grained GUI element detection [hong2023cogagent]. The effectiveness of these models can be further amplified by advanced prompting techniques like Set-of-Mark, which significantly improves visual grounding on crowded interfaces [yang2023set].
Beyond static perception, agents must understand dynamic scenes and dialog. This requires models that can perform efficient temporal filtering for tasks like video grounding and understand context through mechanisms like graph inference and knowledge distillation and multimodal dialogue understanding toolkits [li2023v]. The ability to propagate information along event timelines, as explored in audio-visual research , is also conceptually relevant for tracking state changes in a GUI over time. Furthermore, a comprehensive understanding of on-screen content is often necessary, which is being addressed by dedicated benchmarks for visual question answering on screens [li2024screenvqa] and multimodal dialogue understanding toolkits [li2023v]. The reasoning capabilities of these agents are powered by the underlying LLM architecture, which has shown emergent abilities in planning and decomposition when guided by frameworks like Chain-of-Thought [wei2022chain] and can be extended to perform complex algorithmic reasoning [zelikman2024parsel].
@@ II-B Autonomous Agents and Frameworks
An agent is more than just a perception model; it requires an agentic framework that enables it to plan and execute actions to achieve a goal. The ReAct framework was a landmark contribution, demonstrating how to effectively interleave reasoning steps with actions to create more robust and transparent agents [yao2023react]. This core idea has been extended in various ways, for example, by enabling agents to learn to use external tools and APIs autonomously . The concept of self-improvement has also gained traction, with frameworks like Reflexion allowing agents to verbally reflect on past mistakes to improve future performance [shinn2023reflexion].
These agentic principles have been applied across different digital domains. In web automation, projects like WebVoyager [he2024webvoyager] and others [gur2023real] have created agents capable of navigating complex websites. The scope has recently expanded to mobile devices with AppAgent [yang2023appagent] and Mobile-Agent [fu2024mobile], and to general desktop control with pilot studies on Windows [cheng2024windows] and ambitious projects aiming for universal computer control [cradle2024]. The ultimate vision is a generalist agent that can operate in any environment, a concept being pushed forward in complex domains like Minecraft [yuan2024ghost] and robotics [yuan2024towards]. The knowledge required for such generalist agents can be substantial, leading to the development of open-source, knowledge-based agents that can solve complex tasks by querying external information sources [zhou2024ok]. Some approaches even aim to turn the agent into a ”programmer” that translates human instructions into executable code [li2024octopus], blurring the lines between using and creating software.
@@ II-C Learning Paradigms for Interaction
Training these autonomous agents presents a significant research challenge. A primary methodology is imitation learning, or behavioral cloning, where the agent learns by mimicking expert demonstrations. This has been shown to be effective when large, web-scale datasets of commands are available [li2023web]. However, imitation learning alone can suffer from compounding errors and may not generalize well to situations not seen in the training data. To overcome this, Reinforcement Learning (RL) is often employed, allowing agents to learn through trial and error. Early work applied RL to web navigation in richly-observed environments [jia2019reinforcement], and more recent paradigms like the Decision Transformer have elegantly reframed RL as a sequence modeling problem, making it highly compatible with modern LLM architectures [chen2021decision].
Given the high cost of exploration in real-world applications, hybrid approaches are becoming increasingly popular. These methods often pre-train an agent with imitation learning and then fine-tune it with RL. Innovations in this area include self-improvement through human-in-the-loop interaction [zhang2024self] and specialized fine-tuning strategies like Agent-Tuning to better align LLMs with agentic behaviors [zeng2023agentlm]. A critical and often-overlooked aspect of training is data quality. Real-world datasets are frequently noisy, and developing robust methods for label noise suppression is essential for training reliable models . Broader surveys of LLM-based agents also emphasize data quality and robustness . As these models grow in size, deploying them efficiently becomes a challenge, necessitating research into model compression techniques like multi-objective convex quantization . For applications involving sensitive data, privacy-preserving training methods are required. Federated learning offers a promising solution, and recent work has focused on making it practical for heterogeneous edge networks through strategies like experience-driven model migration .
@@ II-D Sensing and Interaction Beyond the GUI
While our work focuses on visual GUI interaction, the broader field of human-computer interaction is exploring a rich tapestry of sensing modalities that provide valuable context for future generalist agents. For instance, WiFi-based sensing has been used for attention-based gesture recognition [10090421], and this line of work is complementary to advances in screen-content understanding for instruction following [li2024screenvqa]. Robustness to open-set conditions—where unknown gestures may be present—has been investigated by frameworks like WiOpen [10899398] , with complementary insights from general-assistant evaluations , and is related to benchmarking finer-grained behaviors such as micro-actions . WiFi signals can also be used for human activity recognition, with techniques developed to mitigate co-channel interference [9075376], alongside broader evaluations of general AI assistants that stress realistic task settings . The fusion of WiFi and vision further enables unobtrusive emotion recognition through combined analysis of gestures and facial expressions [10149418], which dovetails with efforts toward unified, multi-sensor autonomy [wang2023unisim]. In healthcare, non-contact pulmonary function monitoring with commodity WiFi has been demonstrated , and complementary perspectives on decision-making pipelines emphasize reliability [fu2024foundation]. This target-oriented sensing paradigm aligns with insights from generalist robotics [yuan2024towards]; beyond WiFi, other radio-frequency technologies like RFID can recognize fine-grained activities such as writing [feng2025rf]. However, the reliance on physical-layer information opens security attack surfaces—e.g., undetectable attacks on PHY-layer fingerprinting for WiFi authentication [wang2023unisim]—underscoring the need for rigorous testing [araujo2024can].
@@
The fusion of multiple modalities is a particularly promising frontier; for example, combining WiFi and vision has enabled unobtrusive emotion recognition through joint analysis of gestures and facial expressions [10149418]. This mirrors the architectural principles of powerful vision-language models like Flamingo, which are designed to process and reason about multimodal inputs [10804189, alayrac2022flamingo]. The application of these sensing technologies extends even to personal healthcare, where target-oriented sensing methods demonstrate significant potential . This paradigm of non-contact monitoring is a key application area for future foundation models for decision making [fu2024foundation], with specific systems successfully capturing pulmonary function using commodity WiFi [10699348]. In parallel, unified multi-sensor autonomy initiatives highlight cross-modal deployment challenges [wang2023unisim]. Furthermore, other radio-frequency technologies like RFID are being leveraged to recognize fine-grained activities such as writing [feng2025rf]. This multi-sensor vision for autonomy, pursued in frameworks like UniSim , also introduces new security considerations. The reliance on physical layer information opens up novel attack surfaces, necessitating research into defending against threats like undetectable attacks on WiFi authentication systems [10339891].